OpenStudy (anonymous):

describe the discriminant and the nature of the roots: y=-2x^2-4x+5

OpenStudy (anonymous):

please also explain

OpenStudy (jhannybean):

your equation is in the form : \(ax^2+bx+c=0\) Therefore you can find your roots by using the discriminant, \(b^2-4ac\), testing whether it's greater than, less than or equal to 0.

OpenStudy (jhannybean):

First you want to identify \(a~,~ b ~,~ c\).

OpenStudy (anonymous):

so -24?

OpenStudy (anonymous):

hello?

OpenStudy (jhannybean):

Are you sure it's -24?

OpenStudy (jhannybean):

Can you identify what a, b, and c are for me?

OpenStudy (anonymous):

with -2x being a, -4x being b, and 5 being c yes

OpenStudy (jhannybean):

Alright, then \(b^2-4ac = (-4)^2 -4((-2)(5)) = 16 +40 = 56\)

OpenStudy (jhannybean):

I think what you had done was 16 - 40.

OpenStudy (jhannybean):

Are you following?

OpenStudy (jhannybean):

@mathacomplice ?

OpenStudy (anonymous):

oh sorry yes

OpenStudy (anonymous):

i was getting something to eat

OpenStudy (jhannybean):

So the discriminant tells us if: \(b^2-4ac > 0\) We will have \(\color{red}{\text{2 real solutions}}\) ONLY IF we factor our quadratic and end up with a perfect square. \(b^2-4ac = 0\) and your function is NOT a perfect square, then you will have \(\color{red}{\text{2 real irrational solutions}}\) \(b^2 -4ac < 0\) Your function will have no \(\color{red}{\text{real}}\) solutions, but \(\color{red}{\text{2 imaginary/complex solutions}}\)

OpenStudy (jhannybean):

Our discriminant: \(b^2 -4ac = 56 \implies 56 >0\) Therefore we have 2 real solutions.

OpenStudy (jhannybean):

Understand @mathacomplice ?

OpenStudy (anonymous):

ahh, yes thank you very much

OpenStudy (jhannybean):

No problem :)